OpenStudy (anonymous):

Medals medals! and become a fan! helps? what is true about the complex number 5-5i? a. the modulus us ten b. the modulus is 5sqrt2 c. lies in quadrant 2 d. lies in quadrant 4 more than one answer!

OpenStudy (anonymous):

@Hero @jim_thompson5910 can you walk me through this?

jimthompson5910 (jim_thompson5910):

Are you able to plot this complex number on the xy plane?

OpenStudy (anonymous):

im assuming so since i need the quadrants, but im very shaky at imaginary numbers

jimthompson5910 (jim_thompson5910):

Example: 2 + 3i = a+bi a = 2 b = 3 

jimthompson5910 (jim_thompson5910):



OpenStudy (anonymous):

okay, thats clear!

OpenStudy (anonymous):

so im in the 4th quadrant with my question?